# Towards bandit-based prompt-tuning for in-the-wild foundation agents

### v2

@@ Abstract
Prompting has emerged as the dominant paradigm for adapting large, pre-trained transformer-based models to downstream tasks. The Prompting Decision Transformer (PDT) enables large-scale, multi-task offline reinforcement learning pre-training by leveraging stochastic trajectory prompts to identify the target task. However, these prompts are sampled uniformly from expert demonstrations, overlooking a critical limitation: Not all prompts are equally informative for differentiating between tasks. To address this, we propose an inference time bandit-based prompt-tuning framework that explores and optimizes trajectory prompt selection to enhance task performance. Our experiments indicate not only clear performance gains due to bandit-based prompt-tuning, but also better sample complexity, scalability, and prompt space exploration compared to prompt-tuning baselines.
@@ 1 Introduction
Recent advances in Artificial Intelligence (AI) research have demonstrated the strength of large, pre-trained transformer-based foundation models in many domains, including language [Radford, 2018, Brown et al., 2020], vision [Radford et al., 2021, Dosovitskiy, 2020], and reinforcement learning [Reed et al., , Li et al., 2023]. These large models leverage vast and diverse offline datasets to acquire generalizable representations that can solve many downstream tasks. A prominent strategy for leveraging these models in zero- and few-shot settings involves conditioning them on a prompt – a structured input that specifies the current objective. By keeping the prompt in context, the model ensures that subsequently generated tokens are aligned with the task. Consequently, the performance of a pre-trained model in a downstream task is contingent not only on the coverage of the pre-training data but also on the quality and informativeness of the provided prompt [Hu et al., 2023, Lin et al., 2023, Lester et al., 2021].
@@
Building on the success of transformer-based multi-task language models, Offline Reinforcement Learning (ORL) has increasingly adopted transformer architectures, such as the Decision Transformer (DT) [Chen et al., 2021], to address sequential decision-making problems. In the multi-task setting, DT has been extended to the Prompting Decision Transformer (PDT) [Xu et al., 2022], which leverages stochastic trajectory prompts – multiple segments of expert demonstrations – to enable task-conditioned pre-training and to facilitate few-shot adaptation. These prompts serve as task descriptors that allow PDT to distinguish tasks and to generate actions aligned with the optimal policy distribution for each task. However, PDT samples these prompts uniformly at random from the demonstration dataset, overlooking a crucial limitation: We posit that the informativeness of prompts (and, conversely, the segments that compose them) can vary considerably, which can diminish PDT’s ability to identify the target task and lead to performance degradation.
@@
To address this limitation, we introduce a simple yet effective bandit-based prompt-tuning framework that actively explores the prompt space. By formulating prompt selection as a contextual bandit problem, our method systematically identifies and exploits prompts that maximize downstream task performance, without requiring costly modifications to the pre-trained Transformer backbone. This approach is scalable, computationally efficient, and seamlessly integrates with PDT, enhancing performance while eliminating the need for additional task-specific fine-tuning.
@@
We validate the effectiveness of our approach in a controlled proof-of-concept environment, demonstrating that adaptive prompt selection significantly improves the performance of a pre-trained PDT model. Furthermore, ongoing work explores extending this method to more complex environments. Our findings underscore the critical role of prompt optimization in offline RL and reinforce the broader significance of prompt quality in transformer-based decision-making models.
@@ 2 Preliminaries
This section covers the background of our method. We define our learning objective in Sec. 2.1, formalize the contextual bandit problem in Sec. 2.2, and review the PDT [Xu et al., 2022] architecture in Sec. 2.3.
@@ 2.1 Offline multi-task RL
$J(\pi,\rho)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma_{i}^{t}r_{i}(\mathbf{s}_{t},\mathbf{a}_{t})\right]$ (1)
@@ 2.2 Contextual Multi-Armed Bandits
Multi-Armed Bandits (MABs) provide a framework for optimizing stochastic reward functions over the course of $K$ rounds. For each round $k\in\{1,\dots,K\}$ the bandit selects an action $a_{k}\in\mathcal{A}_{b}$ by pulling one of its arms, where $\mathcal{A}_{b}$ denotes the bandit’s set of arms. It then perceives a stochastic reward $r_{k}\sim R(a_{k})$ for performing that action, where $R$ is the reward distribution. The goal is to maximize the cumulative reward $\sum_{k=1}^{K}r_{k}$ over the $K$ rounds, which requires balancing exploration and exploitation of the available arms while minimizing cumulative regret [Auer et al., 2002]:
@@
Contextual Multi-Armed Bandits (CMABs) extend standard MABs by incorporating additional information (i.e. “context”) $\mathbf{c}_{k}\in\mathcal{C}$ observed at each round $k$. The stochastic reward depends on both the action and the context, $r_{k}\sim R(a_{k}\mid\mathbf{c}_{k})$, meaning a CMAB’s objective is to learn a policy $\pi:\mathcal{C}\to\mathcal{A}_{b}$ that maximizes the expected reward objective $\mathbb{E}[\sum_{k=1}^{K}R(\pi(\mathbf{c}_{k})\mid\mathbf{c}_{k})]$. By exploiting the cross-arm features given by the context, CMABs are credited with better sample efficiency and generalization than their non-contextual counterparts [Li et al., 2010], making them well-suited for efficient prompt-tuning.
$\rho=\big{(}\overbrace{\hat{r}_{j},\mathbf{s}_{j},\mathbf{a}_{j},\dots,\hat{r}_{j+H},\mathbf{s}_{j+H},\mathbf{a}_{j+H}}^{\text{ $\tilde{\tau}_{1}$: segment 1}},\dots,\overbrace{\hat{r}_{k},\mathbf{s}_{k},\mathbf{a}_{k},\dots,\hat{r}_{k+H},\mathbf{s}_{k+H},\mathbf{a}_{k+H}}^{\text{$\tilde{\tau}_{J}$: segment }J}\big{)}$ (3)
$\mathbf{x}=\big{(}\rho\big{)}\odot\big{(}\overbrace{\hat{r}_{t-K},\mathbf{s}_{t-K},\mathbf{a}_{t-K},\hat{r}_{t-K+1},\mathbf{s}_{t-K+1},\mathbf{a}_{t-K+1}\dots,\hat{r}_{t},\mathbf{s}_{t},\mathbf{a}_{t}}^{\tau_{N:t}\>N\ \text{most recent transitions}}\big{)}$ (4)
@@ 2.3 Prompting Decision Transformer
With PDT, Xu et al. [2022] treat offline multi-task RL as a sequence learning problem by autoregressively modeling the trajectories in the available offline datasets. Trajectories consist of $(\hat{r}_{t},\mathbf{s}_{t},\mathbf{a}_{t})$ triplets, with $\hat{r}_{t}=\sum^{T}_{t^{\prime}=t}r_{t^{\prime}}$ being return-to-go, needed for conditioning on optimal return. For all training tasks $\mathcal{T}^{\text{train}}$, PDT learns to model the sequence $\mathbf{x}$ in Eq. (4) by autoregressively predicting the action tokens, where $\odot$ denotes concatenation. The prompt $\rho$ consists of $J$ segments, each of length $H$, which Xu et al. [2022] sample uniformly from the expert demonstrations $\mathcal{P}_{i}$ for that task. Instead of relying on uninformed random sampling, we hypothesize that prompts can vary in their usefulness for describing the downstream task, based on segment composition, and propose to optimize prompt and segment selection with a CMAB approach. As we detail in the next section, the bandit explores directly in the prompt space and learns to select the best prompt constructible from $\mathcal{P}_{i}$.
@@ 3 Method: Prompt-tuning contextual bandit
We propose a contextual multi-armed bandit (CMAB) architecture to optimize the prompt selection and segment composition to improve the performance of a pre-trained PDT backbone on a downstream task $\mathcal{T}_{i}$. To this end, we assume access to a PDT $\theta^{*}$, pre-trained until convergence on a multi-task dataset $\mathcal{D}$, a small number of expert demonstrations $\mathcal{P}_{i}$ to select prompts from, and a simulator $\mathcal{M}_{i}$ for the downstream task $i$.
At a high level, our approach operates as follows. For each round $k\in\{1,\dots,K\}$, the bandit selects a prompt $\rho_{k}$ from $\mathcal{P}_{i}$ which is prepended to the PDT’s input according to Eq. (4). We then proceed by rolling out the PDT, conditioned on $\rho_{k}$, in $\mathcal{M}_{i}$ and take note of the achieved online return $G_{k}=\sum_{t=0}^{T}r_{i}(\mathbf{s}_{t},\mathbf{a}_{t})\mid\mathbf{a}_{t}\sim\pi(\mathbf{x}_{t},\theta^{*})$ for that round. Note that while $\tau_{N:k}$ in Eq. (4) is dynamically updated to reflect the last $N$ steps in the episode, the prompt remains fixed during an entire episode. From the bandit’s perspective, $G_{k}$ serves as a reward for selecting prompt $\rho_{k}$, and the tuple $\langle\rho_{k},G_{k}\rangle$ is stored for training the bandit’s reward model.
We now detail our CMAB architecture and how it constructs prompts at each round $k$. Instead of relying on a naïve bandit that employs one arm per prompt constructible from $\mathcal{P}_{i}$ (which would scale linearly with the size of $\mathcal{P}_{i}$ and combinatorially with the number of segments $J$), we exploit the similarity between prompt segments and employ a contextual bandit, with $J$ arms. Our bandit maintains a separate reward model $\phi_{j}:\tilde{\tau}\to\mathbb{R}$ for each arm $j\in\{1,\dots,J\}$, and treats segments as context. These models estimate the return achieved by PDT $\theta^{*}$ when segment $\tilde{\tau}_{l}$ is placed at position $j$ in the prompt. Thus, at each round $k$, our bandit predicts the reward for each segment in each position, resulting in a prediction matrix $\mathbf{Y}$, with $J$ columns and rows equal to $|\mathcal{P}_{i}|$, the number of segments in the given expert demonstration dataset. To select a prompt, the bandit can either exploit based on accumulated knowledge and $\arg\max\mathbf{Y}$ along the segments’ dimension, or explore using some exploration mechanisms such as $\epsilon$-greedy, or Upper Confidence Bounds (UCB). This approach has high variance because each reward model $\phi_{j}$ assumes that its corresponding segment $\tilde{\tau}_{j}$ is the sole determinant of the observed performance $G_{k}$, thereby ignoring contributions from the other segments in the prompt. Nevertheless, our experiments indicate that this approach works well in practice.
@@ 4 Experiments
Environment: We evaluate the proposed prompt-tuning CMAB in a 2D proof-of-concept environment. This environment features a planar 2D point agent that has to reach a goal coordinate. The state contains the agent’s 2D coordinate at each step $t$. The action space contains two continuous actions for translating on the plane, with the step size being limited by projecting the translation vector on a unit circle with a radius of 0.1. In addition, the action space contains a stop action which allows the agent to terminate the episode. When selected, the episode ends, and the agent receives a sparse reward proportional to its distance from the goal. A bonus of +10 reward is provided (discounted for exceeding the optimal number of steps) for stopping in close proximity of the goal coordinate. To create a multi-task setting, we parameterize tasks by $(r,\alpha)$, the goal’s radius and angle. We discretize the task sparsely using 20 discrete angles $\alpha\in\{0.1\cdot\pi,0.2\cdot\pi,\dots,2\cdot\pi\}$ and three discrete radii $r\in\{0.9,1.9,2.9\}$, yielding a total of 60 tasks.
Offline dataset and pre-training: We collect an offline multi-task dataset by training Proximal Policy Optimization (PPO) [Schulman et al., 2017] for 1M steps on each of the 60 tasks, storing the trajectories as $\mathcal{D}_{i}$. We extract trajectories from the top percentile from $\mathcal{D}_{i}$ to serve as expert demonstrations $\mathcal{P}_{i}$ for that task. We then train PDT, without modifications, on $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{60}\}$ and $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{60}\}$ until convergence; see Xu et al. [2022] for details.

### v3

@@ Abstract
Prompting has emerged as the dominant paradigm for adapting large, pre-trained transformer-based models to downstream tasks. The Prompting Decision Transformer (PDT) enables large-scale, multi-task offline Reinforcement Learning (RL) pre-training by leveraging stochastic trajectory prompts to identify the target task. However, these prompts are sampled uniformly from expert demonstrations, overlooking a critical limitation: not all prompts are equally informative for differentiating between tasks. This limits generalization and adaptation, especially in low-data or open-world settings where sample efficiency is crucial. To address this issue, we propose a lightweight, inference-time, bandit-based prompt-tuning framework. The bandit explores and optimizes trajectory prompt selection to enhance task performance, while avoiding costly fine-tuning of the transformer backbone. Our experiments indicate not only clear performance gains due to bandit-based prompt-tuning, but also better sample complexity, scalability, and prompt space exploration compared to prompt-tuning baselines. These results highlights the importance of adaptive prompt selection mechanisms for efficient generalization in offline multi-task RL.
@@ 1 Introduction
Recent advances in Artificial Intelligence (AI) research have demonstrated the strength of large, pre-trained transformer-based foundation models in many domains, including language [12, 2], vision [13, 5], and reinforcement learning [14, 10]. These large models leverage vast and diverse offline datasets to acquire generalizable representations that can solve many downstream tasks. A prominent strategy for leveraging these models in zero- and few-shot settings involves conditioning them on a prompt – a structured input that specifies the current objective. By keeping the prompt in context, the model ensures that subsequently generated tokens are aligned with the task. Consequently, the performance of a pre-trained model in a downstream task is contingent not only on the coverage of the pre-training data but also on the quality and informativeness of the provided prompt [6, 11, 8].
@@
Building on the success of transformer-based multi-task language models, Offline Reinforcement Learning (ORL) has increasingly adopted transformer architectures, such as the Decision Transformer (DT) [4], to address sequential decision-making problems. In the multi-task setting, DT has been extended to the Prompting Decision Transformer (PDT) [19], which leverages stochastic trajectory prompts, multiple segments of expert demonstrations, to enable task-conditioned pre-training and to facilitate few-shot adaptation. These prompts serve as task descriptors that allow PDT to distinguish tasks and to generate actions aligned with the optimal policy distribution for each task. However, PDT samples these prompts uniformly at random from the demonstration dataset, overlooking a crucial limitation: not all prompts are equally informative for identifying the target task, which can lead to performance degradation and hinder generalization capabilities when uninformed sampling methods are employed. This limitation is especially problematic in applications where online exploration is costly and where adaptation to the target task must happen efficiently and robustly.
Although prompt-tuning is becoming increasingly popular to increase the few-shot performance of Large Language Models (LLMs), e.g. [8, 11, 16, 3], prompt-tuning of RL decision transformers is still underexplored. Hu et al. [6, 7] propose prompt-tuning for PDT by estimating the prompt’s gradient with respect to online task return and by incrementally updating the prompt. However, performing gradient ascent directly on the prompt tokens ignores causal structures in the prompt and is not applicable in discrete settings. Yuan et al. [20] also propose a prompt-tuning approach for PDT, however, their method relies online exploration and additional data collection in the downstream tasks.
@@
Thus, to address the limitations of PDT and prior prompt-tuning works, we introduce a simple yet effective bandit-based prompt-tuning framework that actively explores the prompt space. By formulating prompt selection as a contextual bandit problem, our method systematically identifies and exploits prompts that maximize downstream task performance, without requiring costly modifications to the pre-trained Transformer backbone. This approach is scalable, computationally efficient, and seamlessly integrates with PDT, enhancing performance while eliminating the need for additional task-specific fine-tuning. We summarize our contributions as follows:
  • We introduce a bandit-based prompt-tuning approach to explore the prompt space at inference time, aiming to find the most effective prompt sequence for each downstream task.
@@
  • We validate the effectiveness of our approach in a controlled proof-of-concept environment, demonstrating that adaptive prompt selection significantly improves the performance of a pre-trained PDT model.
  • We benchmark our method against two optimization-based prompt-tuning baselines and evaluate it using two bandit strategies across varying numbers of prompt segments.
  • We provide analysis on exploration effectiveness and sensitivity to the initialization among different approaches.
Furthermore, ongoing work explores extending this method to more complex environments. Our findings underscore the critical role of prompt optimization in offline RL and reinforce the broader significance of prompt quality in transformer-based decision-making models.
@@ 2 Preliminaries
This section covers the background of our method. We define our learning objective in Sec. 2.1, formalize the contextual bandit problem in Sec. 2.2, and review the PDT [19] architecture in Sec. 2.3.
@@ 2.1 Offline multi-task RL
$J(\pi,i)=\mathbb{E}_{\mathbf{a}_{t}\sim\pi(\rho),\mathbf{s_{t}}\sim d_{i}}\left[\sum_{t=0}^{\infty}\gamma_{i}^{t}r_{i}(\mathbf{s}_{t},\mathbf{a}_{t})\right],\rho\sim\mathcal{P}_{i}$ (1)
@@ 2.2 Contextual Multi-Armed Bandits
Multi-Armed Bandits (MABs) provide a framework for optimizing stochastic reward functions over the course of $K$ rounds. For each round $k\in\{1,\dots,K\}$ the bandit selects an action $a_{k}\in\mathcal{A}_{b}$ by pulling one of its arms, where $\mathcal{A}_{b}$ denotes the bandit’s set of arms. It then perceives a stochastic reward $r_{k}\sim R(a_{k})$ for performing that action, where $R$ is the reward distribution. The goal is to maximize the cumulative reward $\sum_{k=1}^{K}r_{k}$ over the $K$ rounds, which requires balancing exploration and exploitation of the available arms while minimizing cumulative regret [1]:
@@
Contextual Multi-Armed Bandits (CMABs) extend standard MABs by incorporating additional information (i.e. “context”) $\mathbf{c}_{k}\in\mathcal{C}$ observed at each round $k$. The stochastic reward depends on both the action and the context, $r_{k}\sim R(a_{k}\mid\mathbf{c}_{k})$, meaning a CMAB’s objective is to learn a policy $\pi:\mathcal{C}\to\mathcal{A}_{b}$ that maximizes the expected reward objective $\mathbb{E}[\sum_{k=1}^{K}R(\pi(\mathbf{c}_{k})\mid\mathbf{c}_{k})]$. By exploiting the cross-arm features given by the context, CMABs are credited with better sample efficiency and generalization than their non-contextual counterparts [9], making them well-suited for efficient prompt-tuning.
@@ 2.3 Prompting Decision Transformer
Prompting Decision Transformer (PDT) [19] treats offline multi-task RL as a sequence learning problem by autoregressively modeling the trajectories in the available offline datasets. Trajectories consist of $(\hat{r}_{t},\mathbf{s}_{t},\mathbf{a}_{t})$ triplets, with $\hat{r}_{t}=\sum^{T}_{t^{\prime}=t}r_{t^{\prime}}$ being return-to-go, needed for conditioning on optimal return. For all training tasks $\mathcal{T}^{\text{train}}$, PDT learns to model the sequence $\mathbf{x}$ in Eq. (4) by autoregressively predicting the action tokens, where $\odot$ denotes concatenation. The prompt $\rho$ in Eq. (3) consists of $J$ segments, each of length $H$, that are sampled uniformly from the expert demonstrations $\mathcal{P}_{i}$ for that task.
$\rho=\big{(}\overbrace{\hat{r}_{l},\mathbf{s}_{l},\mathbf{a}_{l},\dots,\hat{r}_{l+H},\mathbf{s}_{l+H},\mathbf{a}_{l+H}}^{\text{ $\tilde{\tau}_{1}$: segment 1}},\dots,\overbrace{\hat{r}_{k},\mathbf{s}_{k},\mathbf{a}_{k},\dots,\hat{r}_{k+H},\mathbf{s}_{k+H},\mathbf{a}_{k+H}}^{\text{$\tilde{\tau}_{J}$: segment }J}\big{)}$ (3)
$\mathbf{x}=\big{(}\rho\big{)}\odot\big{(}\overbrace{\hat{r}_{t-N},\mathbf{s}_{t-N},\mathbf{a}_{t-N},\hat{r}_{t-N+1},\mathbf{s}_{t-N+1},\mathbf{a}_{t-N+1}\dots,\hat{r}_{t},\mathbf{s}_{t},\mathbf{a}_{t}}^{\tau_{N:t}\>N\ \text{most recent transitions}}\big{)}$ (4)
Instead of relying on uninformed random sampling, we hypothesize that prompts can vary in their usefulness for describing the downstream task, based on segment composition and overlap between demonstrations for multiple tasks, and propose to optimize prompt and segment selection with a CMAB approach. As we detail in the next section, the bandit explores directly in the prompt space and learns to select the best prompt constructible from $\mathcal{P}_{i}$.
@@ 3 Method: Prompt-tuning contextual bandit
We propose a contextual multi-armed bandit (CMAB) architecture to optimize the prompt selection and segment composition to improve the performance of a pre-trained PDT backbone on a downstream task $\mathcal{T}_{i}$. To this end, we assume access to a PDT $\theta^{*}$, pre-trained until convergence on a multi-task dataset $\mathcal{D}$, a small number of expert demonstrations $\mathcal{P}_{i}$ to select prompts from, and a simulator $\mathcal{M}_{i}$ to evaluate online performance for the downstream task $i$.
At a high level, our approach operates as follows. For each round $k\in\{1,\dots,K\}$, the bandit selects a prompt $\rho_{k}$ from $\mathcal{P}_{i}$ which is prepended to the PDT’s input according to Eq. (4). We then proceed by rolling out the PDT, conditioned on $\rho_{k}$, in $\mathcal{M}_{i}$ and take note of the achieved online return $G_{k}=\sum_{t=0}^{T}r_{i}(\mathbf{s}_{t},\mathbf{a}_{t})\mid\mathbf{a}_{t}\sim\pi(\mathbf{x}_{t},\theta^{*})$ for that round. Note that while $\tau_{N:t}$ in Eq. (4) is dynamically updated to reflect the last $N$ steps in the episode, the prompt remains fixed during an entire episode. From the bandit’s perspective, $G_{k}$ serves as a reward for selecting prompt $\rho_{k}$, and the tuple $\langle\rho_{k},G_{k}\rangle$ is stored for training the bandit’s reward model.
We now detail our CMAB architecture and how it constructs prompts at each round $k$. A simple but naïve approach to bandit based prompt-tuning could be to maintain one arm per prompt constructible from the demonstrations in $\mathcal{P}_{i}$, however, the search space would grow combinatorially with the number of segments $J$ in the prompt under this architecture. We instead propose a contextual bandit with only $J$ arms, one for each segment in the prompt, taking the structure of the prompt space and similarity between segments into consideration. Our bandit maintains a separate reward model $\phi_{j}:\tilde{\tau}\to\mathbb{R}$ for each arm $j\in\{1,\dots,J\}$, and treats prompt segments as context. Each of these reward models estimates the return achieved by the PDT for task $i$, when segment $\tilde{\tau}\in\mathcal{P}_{i}$ is placed at position $j$ in the prompt. Thus, at each round $k$, our bandit predicts the reward for each segment in each position of the prompt, resulting in a prediction matrix $\mathbf{Y}$, with $J$ columns and rows equal to $|\mathcal{P}_{i}|$, the number of segments in the given expert demonstration dataset. To select a prompt, the bandit can either exploit based on accumulated knowledge and $\arg\max\mathbf{Y}$ along the segments’ dimension, or explore using some exploration mechanisms such as $\epsilon$-greedy, Upper Confidence Bounds (UCB) [1], or Thompson Sampling [18]. To summarize, we propose a bandit architecture tailored to efficient prompt-tuning of PDT models and maintain linear rather then combinatorial scaling with the prompt size.
@@ 4 Experiments
Environment: We evaluate our proposed prompt-tuning bandit architecture in a 2D proof-of-concept environment. This environment features a planar 2D point agent that has to reach a goal coordinate. The state contains the agent’s 2D coordinate at each step $t$. The action space contains two continuous actions for translating on the plane, with the step size being limited by projecting the translation vector on a unit circle with a radius of 0.1. In addition, the action space contains a binary stop action which allows the agent to terminate the episode.
@@
When selected, the episode ends, and the agent receives a sparse reward proportional to its distance from the goal. A bonus of +10 reward is provided (discounted for exceeding the optimal number of steps) for stopping in close proximity of the goal coordinate. To create a multi-task setting, we parameterize tasks by $(r,\alpha)$, the goal’s radius and angle. We discretize the task sparsely using 20 discrete angles $\alpha\in\{0.1\cdot\pi,0.2\cdot\pi,\dots,2\cdot\pi\}$ and three discrete radii $r\in\{0.9,1.9,2.9\}$, yielding a total of 60 tasks.
Offline dataset and pre-training: We collect an offline multi-task dataset by training Proximal Policy Optimization (PPO) [15] for 1M steps on each of the 60 tasks, storing the trajectories as $\mathcal{D}_{i}$. We extract trajectories from the top percentile from $\mathcal{D}_{i}$ to serve as expert demonstrations $\mathcal{P}_{i}$ for that task. We then train PDT, without modifications, on $\mathcal{D}=\{\mathcal{D}_{1},\mathcal{D}_{2},\dots,\mathcal{D}_{60}\}$ and $\mathcal{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{60}\}$ until convergence; see [19] for details.